A ball of mass 1000kg is dropped from a height of 5m and rebounded to a height of 25cm calculate:(I) it's kinetic energy just before impact (ii) its initial rebounce velocity and kinetic energy

1
Expert's answer
2022-01-26T07:40:04-0500

(i) it's kinetic energy just before impact:


"KE=mgh=49000\text{ J}."

(ii) its initial rebounce velocity and kinetic energy


"v_r=\sqrt{2gh}=2.21\text{ m/s}.\\\space\\
KE_r=\dfrac 12mv_r^2=2442\text{ J}."